OpenStudy (anonymous):

47 °F = ____ °C (give answer in decimal form, to nearest tenth)

OpenStudy (whpalmer4):

do you know the formula?

OpenStudy (whpalmer4):

0 C = 32 F 100 C = 212 F if you have C, to get F, multiply by 9/5 and add 32. If you have F, to get C, subtract 32 and multiply by 5/9

OpenStudy (anonymous):

i dont get it

OpenStudy (whpalmer4):

here's the formula to give you the temperature in degrees C if you have the temperature in degrees F C = (F-32)*5/9 where F is the temperature in degrees F and C is the temperature in degrees C. For your problem: \[C = (F-32)*5/9 = (47-32)*5/9 = \]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

the answer is 8.3333 and keeps on going.

OpenStudy (whpalmer4):

No, for this problem the answer is 8.3, because the problem statement asks for the number to be rounded to the nearest tenth :-)